The HIRS test data should have looked better than the operational
data last week. We discovered that the operational 24 hour calibration
file was not being updated. This was changed late Friday
afternoon. Now, when we compare the calibration information from the ops
and test, we see the same values for calibration. If you can take one
more look at it before we turn on the clock corrections, you should find
the same results in ops and test for the calibration parameters.
We plan to turn on the clock corrections on Wednesday around 11:30 am local
time. The last orbit to be processed before the change is B1104848,WI.
(LAC) the first orbits to be processed after the change is
B1105151.GC(HRPT), B1105051.GC (GAC), and B1104950.GC(GAC)
Thanks for your support.
Emily D. Harrod
E/SP13 Room 0308 FB#4
5200 Auth Road
Suitland, Md. 20746-4304
Cecil A. Paris
Calibration Physical Scientist